If a buffer solution is 0.210 M in a weak acid (K, 1 2 x105) and 0.540 M in its conjugate base, what is the pH? pH
Answer #1

Given, Ka = 1.2 * 10-5

pKa = - Log (1.2 * 10-5)

pKa = 4.92

[Acid] = 0.210 M

[Conjugae base] = 0.540 M

According Henderson's and Hessalbatch's equation of buffers,

pH = pKa + Log[conjugate base] / [acid]

pH = 4.92 + Log(0.540/0.210)

pH = 5.33